OpenStudy (anonymous):

Find the exact solution of the equation. 3arctan(2x) = π

OpenStudy (anonymous):

start with \(\arctan(2x)=\frac{\pi}{3}\)

OpenStudy (anonymous):

then what do i do

OpenStudy (anonymous):

hmm that is a good question

OpenStudy (anonymous):

maybe a good idea is to take the tangent of both sides

OpenStudy (anonymous):

since \(\tan(\arctan(2x))=2x\) you get \(2x=\tan(\frac{\pi}{3})\)

OpenStudy (anonymous):

now compute the \(\tan(\frac{\pi}{3})\) and then divide by 2

OpenStudy (anonymous):

you should get \[2x=\sqrt{3}\] and so \(x=\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

Thank you so much! This was very helpful! :)

OpenStudy (anonymous):

yw